OpenStudy (anonymous):

Otto mixed spices to create his own Italian seasoning. He used %10 garlic powder. Which of the following shows the percent equation and solution for how many tablespoons of garlic powder he uses to create 20 tablespoons of the seasoning?

OpenStudy (anonymous):

20 tablespoons x 10% garlic powder is written as 20 x .1, which gives us 2. Otto made some pretty decent Italian seasoning.

OpenStudy (anonymous):

\[A) 20 = 0.1 \times a; a = 2\] \[B)20 = 0.1 \times a; a = 200\] \[C) a = 0.1 \times 20; a = 200\] \[D)a= 0.1 \times 20; a = 2\]

OpenStudy (anonymous):

Which One?

OpenStudy (anonymous):

need help ?

OpenStudy (anonymous):

and what os the question

OpenStudy (anonymous):

Otto mixed spices to create his own Italian seasoning. He used %10 garlic powder. Which of the following shows the percent equation and solution for how many tablespoons of garlic powder he uses to create 20 tablespoons of the seasoning?